OpenStudy (anonymous):

3√12 in radical form

OpenStudy (anonymous):

First, simply \[\sqrt{12}\]

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

\[\sqrt{12}\]=\[2\sqrt{3}\]

OpenStudy (anonymous):

Oh ty

OpenStudy (anonymous):

Multiply 2x3=6 So you get \[6\sqrt{3}\]

OpenStudy (anonymous):

So the answer is 6\[\sqrt{3}\]

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

No problem. Let me know if you need help. ; ) Gave a medal.